HUBBLE SPACE TELESCOPE DAILY REPORT #1461 PERIOD COVERED: 0000Z (UTC) 08/23/95 - 0000Z (UTC) 08/24/95 Daily Status Report as of 236/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6186 (UV Throughput) The WF/PC-2 was used to make ultraviolet photometric observations of standard star GRW+70D582 in order to measure the ultraviolet throughput of the system. This was done shortly before and after a DECON through all the UV filters in each chip. The observations were completed as planned, with the only reported anomaly being the loss of lock at 235/024533Z reported in yesterday's daily report and in HSTAR 5256. 1.2 Completed Two Sets of FOS/BL 6166 (FOS Cycle 5: Location of Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make wavelength calibration images to measure the stability of the FOS Blue Side wavelength scale and spectral locations. A Taled calibration observations was also taken. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Two Sets of WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration images. Both photometric and narrow-band filters were used. The observations were executed as planned, and no problems were reported. 1.4 Completed FOS/RD 5607 (Probing Excitation Mechanisms and Abundances in Young Magellanic Cloud) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of supernova remnant N132D-P1 which is located in the Large Magellanic Cloud. This is to obtain spectra of three individual knots in the remnant to test emission mechanisms and derive elemental abundances. The observations were completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrent with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 5957 (Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make observations for radio galaxies 3C109.0 and 3C465.0 to study the morphology of these radio galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed GHRS/1 5879 (The Deuterium to Hydrogen Ration in the Interstellar Medium) The High Resolution Spectrograph (Side One) was used to make spectrographic observations of late-type star HD11443 in order to study the interstellar medium. The objective is to determine the deuterium to hydrogen ratio in the interstellar medium in the line of sight to this star. The observations were executed as scheduled, and no anomalies were noted. 1.8 Competed WF/PC-2 5982 (Imaging of a Complete Sample of the Nearest Infrared Quasars) The WF/PC-2 was used to make two color observations of the nearby infrared quasar (IR QSO) IRAS05189. It is believed that this IR QSO resulted from the merger of two gas-rich galaxies. The observations were executed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 At 236/022759Z, one acquisition failed to coarse track. The event was described in HSTAR 5257. Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: There were no reported operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-